OpenStudy (anonymous):

4cos theta +1 = 2cos theta

OpenStudy (anonymous):

\[1= 2 \cos \theta- 4 \cos \theta=-2 \cos \theta, \cos \theta=\frac{ -1 }{ 2 }\] fond theta